Given the constant elasticity demand function as :

𝑃=π‘Žπ‘ƒπ‘ π‘€β„Žπ‘’π‘Ÿπ‘’ 𝑏 𝑖𝑠 π‘‘β„Žπ‘’ 𝑒𝑙𝑠𝑑𝑖𝑐𝑖𝑑𝑦 π‘œπ‘“ π‘‘π‘’π‘šπ‘Žπ‘›π‘‘

a. Show that Marginal Revenue of this function is proportional to the price

(let 𝐾=(1π‘Ž)1𝑏⁄ ). To simplify the equation.

b. calculate the Marginal Revenue when π‘’π‘žπ‘’π‘–π‘™π‘ƒ=𝑏=βˆ’2 and when b=-10.

c. what does your answer mean in terms of revenue facing the firm?

d. If 𝑏=βˆ’π›Ό what would this imply for Marginal Revenue and Total Revenue of the firm.

e. Explain how Marginal Revenue and profit maximization would be affected if demand was inelastic.


1
Expert's answer
2021-10-12T09:54:26-0400

constant utility demand function is given as :

"p=ap^b"

b- elasticity of demand

a.

"p=ap^b\\"


"\frac{1}{a}=\frac{p^b}{p}\\"


"\frac{1}{a}=pb^{-1}"


"p=(\frac{1}{a})^{\frac{1}{b-1}}"


"Revenue=p.Q\\=(\frac{1}{a})^{\frac{1}{b-1}}"


"MR=p"


Therefore, marginal revenue of the function is proportional to the price.


b

when equil"p=b=-2"

"p=ap^b\\(-2)=a(-2)^{-2}"

"a=-8"


"\implies MR=(\frac{1}{-8})^{\frac{1}{-2-1}}\\MR=-2"


when equil "p=b=-10"

"p=ap^b\\(-10)=a(-10)^{-10}\\a=-10^{11}"


"\implies MR =-10"


c.

The marginal revenue, MR is negative in both the cases. this means that the toital revenue facing the firm is decreasing.

d.

"\frac{g}{b} =-\alpha," The marginal revenue of the firm would be "-\alpha" and the total revenue of the firm would be decreasing.